OpenStudy (anonymous):

If the length of a rectangle is increased by 20% and the width is increased by 30%, then by what percent is the area of the rectangle increased?

OpenStudy (blacksteel):

Suppose a rectangle's width is denoted by W and its length is denoted by L. Then its area is A = L*W Then if our rectangle's dimensions change to 1.2*L and 1.3*W, then our area becomes 1.2L*1.3W = 1.56*L*W, which is 56% larger than our original area.

OpenStudy (anonymous):

@Blacksteel thank you :-)